# UP-VLA: A Unified Understanding and Prediction Model for Embodied Agent

### v2

@@ Abstract
Recent advancements in Vision-Language-Action (VLA) models have leveraged pre-trained Vision-Language Models (VLMs) to improve the generalization capabilities. VLMs, typically pre-trained on vision-language understanding tasks, provide rich semantic knowledge and reasoning abilities. However, prior research has shown that VLMs often focus on high-level semantic content and neglect low-level features, limiting their ability to capture detailed spatial information and understand physical dynamics. These aspects, which are crucial for robotic control tasks, remain underexplored in existing pre-training paradigms. In this paper, we investigate the training paradigm for VLAs, and introduce UP-VLA, a Unified VLA model training with both multi-modal Understanding and future Prediction objectives, enhancing both high-level semantic comprehension and low-level spatial understanding. Experimental results show that UP-VLA achieves a 33% improvement on the Calvin ABC-D benchmark compared to the previous state-of-the-art method. Additionally, UP-VLA demonstrates improved success rates in real-world manipulation tasks, particularly those requiring precise spatial information.
@@ 1 Introduction
Constructing Vision-Language-Action (VLA) models (Brohan et al., 2023; Li et al., 2023b) capable of solving multiple tasks in open environments has become a central focus of research in robotics. A promising approach for VLA models involves fine-tuning large-scale pre-trained Vision-Language Models (VLMs) (Li et al., 2023a; Wang et al., 2022; Dai et al., 2024; Driess et al., 2023) on robotic action datasets, incorporating appropriate action modeling components (Jang et al., 2022; Li et al., 2023b; Wu et al., 2023; Zhang et al., 2024; Kim et al., 2024; Zheng et al., 2024b). This method enables VLA models to inherit the semantic knowledge and reasoning capabilities encoded in powerful VLMs, thereby enhancing decision-making in unknown environments.
@@ 2 Related Works
VLA Models for Generalist Robot Policies Recent studies have explored the application of VLMs (Li et al., 2023a; Wang et al., 2022; Dai et al., 2024; Driess et al., 2023; Wang et al., 2023) in robotics, leveraging their strong understanding of linguistic instructions and visual scenes. A notable example is RT-2 (Brohan et al., 2023), which directly utilizes VLMs to generate discrete action tokens autoregressively, demonstrating the semantic grounding capabilities of VLA methods. Recent works aim to enhance VLA models with better generalization performance (Kim et al., 2024; O’Neill et al., 2023), cross-embodiment control capabilities (Black et al., 2024), and improved reasoning efficiency (Zhang et al., 2024). The previous work, 3D-VLA (Zhen et al., 2024), also explored co-training for multi-modal understanding and generation, but focused on introducing 3d information and employed a separate diffusion model for generation. In contrast, our approach uses a unified model to handle multi-modal input and mainly focuses on addressing the limitations of VLA models in visual perception and physical dynamics.
Visual Pretraining Methods for Robotics Leveraging pretrained vision models for robotic perception has become a crucial area of research in robot control. Early works (Brohan et al., 2022; Jang et al., 2022) employed pretrained vision encoders like ViT (Dosovitskiy et al., 2020) and EfficientNet (Tan & Le, 2019) to encode visual observations. Recently, numerous studies have incorporated generative models (Ho et al., 2020; Blattmann et al., 2023) for training policies via future frame prediction (Guo et al., 2024) and video generation (Du et al., 2024). For example, SuSIE (Black et al., 2023) learns robot actions by predicting keyframes, while GR-1 (Wu et al., 2023) directly pretrains policies through video generation. PAD (Guo et al., 2024) employs diffusion models to predict future images and multi-step actions simultaneously. IGOR (Chen et al., 2024b) uses latent actions that compress visual changes as the intermediate goal for low-level actions. These studies highlight that visual prediction tasks can benefit models’ visual generalization to unseen scenes. Our approach leverages autoregressive VLMs to predict future images, capturing physical dynamics with rich visual information.
@@ 5.1 Experiment Setup and baseline
Our real-world experiments involve multiple table-top manipulation tasks on the Franka-Emika Panda robot, including picking and placing, routing cables, pressing buttons, and opening drawers. Specifically, we collect over 2k demonstrations above 6 skills. As shown in Figure 5(b), we train UP-VLA on simple scenes while testing it on more complex settings. We place several seen and unseen objects on the table to introduce distractions and test whether the model can grasp entirely new objects to verify its semantic grounding capabilities. Meanwhile, we evaluate the model’s ability to perform more fine-grained tasks, such as routing cables, grasping smaller unseen blocks, or picking up a pen. More details of dataset can be found in Appendix LABEL:sec:app_data.

### v3

@@ Abstract
Recent advancements in Vision-Language-Action (VLA) models have leveraged pre-trained Vision-Language Models (VLMs) to improve the generalization capabilities. VLMs, typically pre-trained on vision-language understanding tasks, provide rich semantic knowledge and reasoning abilities. However, prior research has shown that VLMs often focus on high-level semantic content and neglect low-level features, limiting their ability to capture detailed visual and spatial information. These aspects, which are crucial for robotic control tasks, remain underexplored in existing pre-training paradigms. In this paper, we investigate the training paradigm for VLAs, and introduce UP-VLA, a Unified VLA model training with both multi-modal Understanding and future Prediction objectives, enhancing both high-level semantic comprehension and low-level spatial understanding. Experimental results show that UP-VLA achieves a 33% improvement on the Calvin ABC-D benchmark compared to the previous state-of-the-art method. Additionally, UP-VLA demonstrates improved success rates in real-world manipulation tasks, particularly those requiring precise spatial information. Code can be found at https://github.com/CladernyJorn/UP-VLA.
@@ 1 Introduction
Constructing Vision-Language-Action (VLA) models (Brohan et al., 2023; Li et al., 2023b) capable of solving multiple tasks in open environments has become a central focus of research in robotics. A promising approach for VLA models involves fine-tuning large-scale pre-trained Vision-Language Models (VLMs) (Li et al., 2023a; Wang et al., 2022; Dai et al., 2024; Driess et al., 2023) on robotic action datasets, incorporating appropriate action modeling components (Jang et al., 2022; Li et al., 2023b; Wu et al., 2023; Zhang et al., 2024; Kim et al., 2024; Zheng et al., 2024b; Cui et al., 2025). This method enables VLA models to inherit the semantic knowledge and reasoning capabilities encoded in powerful VLMs, thereby enhancing decision-making in unknown environments.
@@ 2 Related Works
VLA Models for Generalist Robot Policies Recent studies have explored the application of VLMs (Li et al., 2023a; Wang et al., 2022; Dai et al., 2024; Driess et al., 2023; Wang et al., 2023; Zhao et al., 2025; Ding et al., 2025; Song et al., 2025; Ding et al., 2024) in robotics, leveraging their strong understanding of linguistic instructions and visual scenes. A notable example is RT-2 (Brohan et al., 2023), which directly utilizes VLMs to generate discrete action tokens autoregressively, demonstrating the semantic grounding capabilities of VLA methods. Recent works aim to enhance VLA models with better generalization performance (Kim et al., 2024; O’Neill et al., 2023), cross-embodiment control capabilities (Black et al., 2024), and improved reasoning efficiency (Zhang et al., 2024). The previous work, 3D-VLA (Zhen et al., 2024), also explored co-training for multi-modal understanding and generation, but focused on introducing 3d information and employed a separate diffusion model for generation. In contrast, our approach uses a unified model to handle multi-modal input and mainly focuses on addressing the limitations of VLA models in visual perception and physical dynamics.
Visual Pretraining Methods for Robotics Leveraging pretrained vision models for robotic perception has become a crucial area of research in robot control. Early works (Brohan et al., 2022; Jang et al., 2022) employed pretrained vision encoders like ViT (Dosovitskiy et al., 2020) and EfficientNet (Tan & Le, 2019) to encode visual observations. Recently, numerous studies have incorporated generative models (Ho et al., 2020; Blattmann et al., 2023) for training policies via future frame prediction (Guo et al., 2024; Ding et al., 2024) and video generation (Du et al., 2024). For example, SuSIE (Black et al., 2023) learns robot actions by predicting keyframes, while GR-1 (Wu et al., 2023) directly pretrains policies through video generation. PAD (Guo et al., 2024) employs diffusion models to predict future images and multi-step actions simultaneously. IGOR (Chen et al., 2024b) uses latent actions that compress visual changes as the intermediate goal for low-level actions. These studies highlight that visual prediction tasks can benefit models’ visual generalization to unseen scenes. Our approach leverages autoregressive VLMs to predict future images, capturing physical dynamics with rich visual information.
@@ 5.1 Experiment Setup and baseline
Our real-world experiments involve multiple table-top manipulation tasks on the Franka-Emika Panda robot, including picking and placing, routing cables, pressing buttons, and opening drawers. Specifically, we collect over 2k demonstrations above 6 skills. As shown in Figure 5(b), we train UP-VLA on simple scenes while testing it on more complex settings. We place several seen and unseen objects on the table to introduce distractions and test whether the model can grasp entirely new objects to verify its semantic grounding capabilities. Meanwhile, we evaluate the model’s ability to perform more fine-grained tasks, such as routing cables, grasping smaller unseen blocks, or picking up a pen. More details of dataset can be found in Appendix B.
@@ 6 Conclusion
Impact Statement
This paper presents a novel research to advance robot manipulation models with unifed training strategies. Given that robots operate in the physical world under certain human instructions, high-level semantic content and also low-level visual and spatial details are crucial for accurate robot control. To mitigate this issue, our approach involves unified training paradigm to force VLA to capture both semantic information and learn dynamics of the physical world.
Appendix A Implementation Details
We use pretrained Showo-512x512(1.3B) (Xie et al., 2024) as the backbone and CLIP-VIT (Radford et al., 2021), MagVIT (Yu et al., 2023) (VQ-GAN (Esser et al., 2021)). In the pretrain stage, we train UP-VLA for 20k steps with batch size of 64 on future prediction and vision-language understanding tasks. We apply a linear warmup at the first 1k steps. In the action learning stage, we train UP-VLA with a batch size of 64.
Appendix B Manipulation Dataset Details
For the simulation environment data, we follow the setups of the CALVIN benchmark (using its training sets and evaluation sets).
In the real world, our training data is collected using both manual and scripted methods. Specifically, for tasks such as grasping two types of toy fruits (carrot and eggplant), opening drawers, and routing cable tasks, we collect demonstrations manually using a remote operation joystick, ensuring that the target objects are roughly evenly distributed in the field of view. For grasping blocks of different colors, we use scripted policies. We randomly initialize the robotic arm’s position to collect trajectories for these tasks.
For real-world evaluation, our tests primarily focus on unseen settings to validate the semantic generalization capability of our method. We test whether the model can complete tasks despite the introduction of more distracting objects, a broader range of positional variations, diverse backgrounds, and entirely new objects, e.g. different shapes of vegetables, an arrow-shaped paper, unseen vegetables, toy pizza, and blocks with unseen color. We also design precise tasks to better test model’s ability to handle visual details, like picking up a small ball or pen.